OpenStudy (anonymous):

Choose the pair of numbers that is not a solution to the given equation. y = 5 - 2x My options are; 1,3 2,1 0,3

OpenStudy (anonymous):

Plug each answer into the equation til you find the one that is not true: 3=5-2(1) 3=3 1=5-2(2) 1=1 3=5-2(0) 3=5 The last one is not true, so that is your answer.

OpenStudy (anonymous):

Thank you! Are you willing to message me & help me with more?

OpenStudy (anonymous):

I have my own homework to do today, but I could help you with a few more :)

OpenStudy (anonymous):

For the equation 3x - 5y = -3, what is the value of y when x is 1?

OpenStudy (anonymous):

x=1, so put 1 in place of x in the equation: 3(1)-5y=-3 3-5y=-3 3-3-5y=-3-3 -5y=-6 -5y/-5=-6/-5 y=6/5

OpenStudy (anonymous):

Which of the following equations is a constant function? y=x y=5 x=5

OpenStudy (anonymous):

I think it's y=x

OpenStudy (anonymous):

If y = 2x - 3, then which of the following ordered pairs lies on the graph?

OpenStudy (anonymous):

Actually a constant function is a horizontal line, so that last one is actually y=5

OpenStudy (anonymous):

What are your options for the next one?

OpenStudy (anonymous):

If y = 2x - 3, then which of the following ordered pairs lies on the graph? (1, -1) (-3, 0) (5, 4)

OpenStudy (anonymous):

Plug your points in to see which one is true (meaning it lies on the graph): -1=2(1)-3 -1=2-3 -1=-1 So (1, -1) is your answer.

OpenStudy (anonymous):

Which of the following does not lie on the graph of y = 2x - 2? (-1/2, -3) (1/2, 1) (-1, -4)

OpenStudy (anonymous):

-3=2(-1/2)-2 -3=-1-2 -3=-3 1=2(1/2)-2 1=1-2 1=-1 Answer is (1/2, 1)

OpenStudy (anonymous):

Second and third ones.

OpenStudy (anonymous):

Sorry, I have to go. Hopefully you can take it from here. If not, start a new question and someone else will help you :) Good luck!